1. the recommended daily calorie intake for teenage girls is 2,200 calories/day. a nutritionist at a state university believes the average daily caloric intake of girls in that state to be lower. test that hypothesis, at the 5% level of significance, against the null hypothesis that the population average is 2,200 calories/day using the following sample data: n=36,x¯=2,150,s=203.

1
Expert's answer
2022-06-13T14:58:30-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge2200"

"H_1:\mu<2200"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=35" and the critical value for a left-tailed test is "t_c =- 2.030108."

The rejection region for this left-tailed test is "R = \{t:t<- 2.030108\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{2150-2200}{203/\sqrt{36}}=-1.4778"


Since it is observed that "t=-1.4778>- 2.030108=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for left-tailed, "df=35" degrees of freedom, "t=-1.4778" is "p=0.074202," and since "p=0.074202>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is less than 2200, at the "\alpha = 0.05" significance level.